OpenStudy (anonymous):

What is the slope of the line through the points (2, 6) and (3, 5)? m = _?_

OpenStudy (anonymous):

from 2 to 3 is right one unit from 6 to 5 is down one unit right one, down one, slope is \(-1\)

OpenStudy (anonymous):

you can always compute \(\frac{5-6}{3-2}=\frac{-1}{1}=-1\) but you are trying to answer the question "right one, up or down how many?" so it is easier to do it in your head in this case

OpenStudy (anonymous):

you got this?